OpenStudy (anonymous):

Can someone please help me how to do this x-5 Over X squared - 10x+25

OpenStudy (campbell_st):

\[\frac{x -5}{x^2 - 10x + 25} \] can you factor the denominator...?

OpenStudy (anonymous):

i don't know how to

OpenStudy (anonymous):

First try to find factors of bottom use Ac method for this

OpenStudy (campbell_st):

ok... well that will be a major problem for you in this question.... I'm sure someone will give you the answer....

OpenStudy (anonymous):

Are you familiar with AC method ???

OpenStudy (campbell_st):

I'm surprised that that you are doing a question of this type without knowing how to factor a quadratic...?

OpenStudy (anonymous):

no i don't recall

OpenStudy (campbell_st):

ok.... can you think of 2 factors of 25 that add to -10...?

OpenStudy (anonymous):

yes -5 and -5

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

Now rewrite -10x in terms of -5 and -5 can you try ???

OpenStudy (campbell_st):

thats great and the factors of x^2 are x and x so you problem in factored form is \[\frac{(x -5)}{(x -5)(x -5)}\] remove the common factor... whats left

OpenStudy (anonymous):

hmm so would it be x-5

OpenStudy (anonymous):

?

OpenStudy (campbell_st):

almost... what would be in the numerator..?

OpenStudy (campbell_st):

you have the denominator correct...

OpenStudy (anonymous):

would it be 1?

OpenStudy (campbell_st):

great so the simplified version is \[\frac{1}{x - }\]

OpenStudy (campbell_st):

oops \[\frac{1}{x - 5}\]

OpenStudy (anonymous):

We have to cancel out common terms from top and bottom like  So we get ---- ?

OpenStudy (anonymous):

So we remain with ------ ???

OpenStudy (anonymous):

1 over x-5

OpenStudy (campbell_st):

thats it... well done

OpenStudy (anonymous):

Awesome

OpenStudy (anonymous):

Is it clear to you ????

OpenStudy (anonymous):

thank you all!! you guys explained it very well

OpenStudy (anonymous):

You are always welcome :)

OpenStudy (anonymous):

i really appreciate the help yeah its clear thank you

OpenStudy (anonymous):

would you mind helping me with another one ?

OpenStudy (anonymous):

sure please

OpenStudy (anonymous):

Ask your question ???

OpenStudy (anonymous):

i don't know how to type it in correctly

OpenStudy (anonymous):

you can draw it using draw tool

OpenStudy (anonymous):

its 1 over 5x -2= x-4 over 5x +3

OpenStudy (anonymous):